# AirCNN via Reconfigurable Intelligent Surfaces: Architecture Design and Implementation

### v1

@@ I Introduction
Beyond wireless communication, RISs have also been proposed for analog computation [5, 6, 7]. By adjusting the RIS phase shifts, an edge server can minimize model parameter aggregation errors. Furthermore, the massive array of reflecting elements within an RIS can be regarded as trainable neurons in the layer of a neural network. Thus, RIS technology holds great promise for the realization of physical neural networks, offering faster computation and lower latency. Despite this potential, research on RIS-assisted physical neural networks remains limited [8, 9, 10, 11, 12, 13]. Early work [8] and [9] developed multi-layer RIS structures devised for controlled laboratory environments for performing deep learning tasks. Work in [10] and [11] proposed RISs that enable edge inference by modeling the RIS-programmable wireless channel as hidden over-the-air (OTA) artificial neural network layers. In [12] and [13], the authors utilized RISs to program channel impulse responses, achieving one-dimensional (1D) and two-dimensional (2D) convolutional neural networks (CNNs), respectively. However, these works have not fully exploited the spatial degrees of freedom (DoFs) available for neural network emulation.
In this paper, we propose AirCNN, a novel OTA computation framework that emulates 2D CNNs through joint optimization of multi-RIS phase shifts, transmitter precoders, and receiver combiners. Unlike previous works [8, 9, 10, 11, 12, 13], existing methods cannot be directly extended to realize 2D CNNs. Emulating 2D CNNs poses additional challenges, including incompatibility with traditional data architectures, transceiver design constraints, and transmission protocol limitations. To address these challenges, we investigate two types of 2D CNNs: the classic 2D convolution (Conv2d) and depthwise separable convolution (ConvSD). We propose corresponding transmission architectures and protocols for both RIS-assisted multiple-input single-output (MISO) and RIS-assisted multiple-input multiple-output (MIMO) systems. A comprehensive comparison between these two RIS-aided systems for physical CNN implementation is conducted, focusing on both performance and implementation overhead. Simulation results demonstrate that the proposed methods achieve satisfactory classification accuracy. Moreover, it is shown that multi-RIS setups significantly outperform single-RIS configurations, particularly in line-of-sight (LoS)-dominated wireless channels.
@@ II System model
Fig. 1(a) depicts a conventional CNN architecture, which comprises three modules: initial layers, a middle layer, and final layers. The initial and final layers may contain a large number of neural network layers depending on the specific application, while the middle layer corresponds to the CNN. This paper aims to emulate the CNN layer using wireless hardware to achieve a similar function as shown in Fig. 1(b). In the analog-based CNN architecture, the middle layer consists of precoders at the transmitter, multiple RISs deployed over the air, and combiners at the receiver. We assume that the transmitter and receiver are equipped with $N_{\rm t}$ and $N_{\rm r}$ antennas, respectively. Furthermore, $L$ RISs are deployed, each comprising $M/L$ reflecting elements, where $M$ is the total number of reflecting elements. Let ${{\bf{\bar{H}}}}_{i}$} and ${{\bf{\hat{H}}}}_{i}$} denote the complex equivalent baseband channel matrices from the transmitter to RIS $i$ and from RIS $i$ to the receiver, respectively. The phase-shift matrix of RIS $i$ is denoted by ${{{\bf{\Theta}}_{i}}}={\rm{diag}}\left({{e^{j{\theta_{i,1}}}},{e^{j{\theta_{i,2}}}},\ldots,{e^{j{\theta_{i,M/L}}}}}\right)$, with $\theta_{i,m}$ denoting the $m$-th phase shift.
@@
A key challenge in AirCNN is the joint design of the transmit precoder, receive combiner, and RIS phase-shift matrices to accurately emulate a given digital convolution kernel $\bf\bar{W}$. Here, we consider a general case where multiple kernels are considered. Thus, $\bf\bar{W}$ is a matrix, where each row represents one kernel. This imitation problem can be formulated as
@@
$\qquad\left|{{{\bf{\Theta}}_{i,i}}}\right|=1,{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}i=1,\ldots,M/L,$ (1c)
where ${\bf F}_{1}$ and ${\bf F}_{2}$ denote the transmit precoder and receive combiner, respectively, $P_{\max}$ is the transmit power budget, and $\bf n$ denotes additive white Gaussian noise, following ${\mathbf{n}}\sim{\cal CN}\left({{\bf 0},{\sigma^{2}}{\bf I}}\right)$. The end-to-end channel matrix $\bf H$ is modeled as ${\bf{H}}=\sum\limits_{i=1}^{L}{{{{\bf{\hat{H}}}}_{i}}{{\bf{\Theta}}_{i}}{{{\bf{\bar{H}}}}_{i}}}$. Instead of solving problem (1) via conventional convex optimization techniques, the approach proposed in this paper is to optimize ${\bf F}_{1}$, ${\bf F}_{2}$, and $\bf{\Theta}$} through end-to-end training based on the given loss function.
III Conv2d-based Physical Neural Network
@@ III-A Conv2d MISO
where ${{\mathbf{F}}_{1,i,t}}\in{{\mathbb{C}}^{{N_{{\text{k}}}^{2}}\times{N_{{\text{k}}}^{2}}}}$ and $f_{2,i,t}$$\in{\mathbb{C}}$ denote the $i$-th OFDM carrier-based precoder at the transmitter and the amplification coefficient at the receiver, respectively. In addition, ${{\mathbf{h}}_{t}^{H}}=\sum\limits_{l=1}^{L}{{\mathbf{\hat{h}}}_{l}^{H}{{\mathbf{\Theta}}_{l,t}}{{{\mathbf{\bar{H}}}}_{l}}}$, where ${\bf{\Theta}}_{l,i}$} represents the $l$th RIS phase shift matrix at time slot $t$.
@@ III-B Conv2d MIMO
The differences between Conv2d MISO and Conv2d MIMO are summarized in Table I. In the table, $T_{\rm s}$ denotes the number of transmission slots, $T_{\rm r}$ the number of RIS adjustments, $T_{\rm o}$ the number of OFDM carriers, $T_{\rm p}$ the number of precoder adjustments, and $T_{\rm c}$ the number of combiner adjustments.
@@
IV ConvSD-based Physical Neural Network
@@ IV-B ConvSD MIMO
which is similar to (4), except that only a single precoder and a single combiner are required in this case. Then, the receiver applies $C_{\rm out}$ filters, each with $C_{\rm in}$ kernels of size $1\times 1$, to imitate the pointwise convolution process. We have
@@
A comparison between the parameters of ConvSD MISO and ConvSD MIMO architectures is presented in Table II.
Figure 3: Multi-RIS aided OTA transmission neural network architecture.
@@ V Numerical results
In this section, we present numerical results to evaluate the image classification accuracy achieved by the proposed schemes, based on the Fashion MNIST dataset. The Fashion MNIST consists of 70,000 images with a resolution of $28\times 28$ pixels, categorized into 10 classes. Among them, 60,000 images are used for training, and 10,000 are utilized for testing. The overall multi-RIS-aided OTA transmission neural network architecture is depicted in Fig. 3, which consists of one convolutional (Conv) layer, one real-to-complex (R2C) layer, two batch normalization (BN) layers, two complex ReLU activation layers, two max pool layers, two complex fully connected (FC) layers, one dropout layer, one complex-to-real (C2R) layer, one real FC layer, and a transceiver module. In addition, the Rician fading channel with Rician factor $K$ is considered for both MIMO and MISO systems. Unless otherwise specified, we set $C_{\rm in}=32$, $C_{\rm out}=64$, $N_{\rm wi}=N_{\rm hi}=14$, $N_{\rm k}=3$, $K=3~{\rm dB}$, $N_{t}=9$, $N_{r}=32$ for ConvSD MIMO, $N_{r}=64$ for Conv2d MIMO, $P_{\rm max}=10~{\rm dB}$, and $\sigma^{2}=1$.
@@
In Fig. 5, we study Rician factor $K$ versus classification accuracy for $L=1$, $P_{\rm max}=10~{\rm dB}$, and $M=50$. As $K$ increases, the channel gain improves, resulting in less signal distortion and enhanced classification accuracy. Furthermore, when $K$ is below $10~{\rm dB}$, the ConvSD MISO scheme outperforms the ConvSD MIMO scheme, however, as $K$ increases, ConvSD MIMO scheme eventually surpasses the ConvSD MISO scheme. This behavior is consistent with the explanation provided for Fig. 4. It is also noteworthy that further increase in $K$ does not always lead to continuous improvements in classification accuracy. Beyond a certain threshold, increasing $K$ may even degrade the performance, as will be discussed later in Fig. 7.
@@
In Fig. 7, we study the classification accuracy of the ConvSD MIMO scheme versus $K$ for $L=1$, $L=2$, and $L=5$ under $P_{\rm max}=10~{\rm dB}$ and $M=50$. It can be observed that for $L=1$, the classification accuracy initially increases but eventually decreases as $K$ continues to grow. This can be explained as follows. When $K$ is below $15~{\rm dB}$, the channel is dominated by non-line-of-sight (NLoS), resulting in a high channel rank and abundant DoFs for neural network modification. Thus, increasing $K$ improves classification accuracy by enhancing the channel gain. However, when $K\geq 15~{\rm dB}$, the channel becomes dominated by LoS components. Although the channel gain remains high, the channel rank tends to decrease toward one, limiting the available DoFs, and thereby, degrading the performance. Moreover, we observe that for larger values of $L$, the system performance remains robust even as $K$ increases. This is because a larger $L$ yields a higher effective channel rank, enhancing the DoFs available for end-to-end training.
@@ VI Conclusion
In this paper, we studied RIS-aided MISO and MIMO systems for engineering the ambient wireless channel to implement CNNs via OTA computation. By jointly training the precoder, combiner, and RIS phase-shift matrices, the digital convolutional operation can be effectively emulated using physical neural networks. We investigated two types of CNNs, namely Conv2d and ConvSD, and proposed two transceiver architectures: RIS-aided MISO and RIS-aided MIMO. A comprehensive comparison between these two architectures for physical CNN implementation was conducted, highlighting the trade-offs between performance gains and implementation overhead. Simulation results demonstrated that the proposed architectures can achieve satisfactory classification accuracy while enabling OTA computation. Furthermore, it was shown that multi-RIS deployments significantly outperform single-RIS when LoS propagation dominates the wireless channel.

### v2

# CNNs in the Air via Reconfigurable Intelligent Surfaces

@@ I Introduction
Beyond wireless communication, RISs have also been proposed for computation [11, 7, 1]. By adjusting the RIS phase shifts, an edge server can minimize model parameter aggregation errors. Furthermore, the massive array of reflecting elements within an RIS can be regarded as trainable neurons in a neural network, thereby enabling tight computation–communication co-design. As a result, RIS technology holds great promise for realizing wireless physical neural networks (WPNNs) with significantly reduced latency and faster inference [5]. Despite this potential, research on RIS-assisted WPNNs remains limited [9, 10, 13, 4, 14]. Early work in [9] developed multi-layer RIS structures devised for controlled laboratory environments for performing deep learning tasks. Work in [10] and [13] proposed RISs that enable edge inference by modeling the RIS-programmable wireless channel as hidden over-the-air (OTA) artificial neural network layers. In [4] and [14], the authors utilized RISs to program channel impulse responses, achieving one-dimensional (1D) and two-dimensional (2D) convolutional neural networks (CNNs), respectively. Although these studies provide important insights into OTA neural computation, extending such designs to CNNs and multi-antenna systems is nontrivial and has been largely unexplored. Particularly, mapping such multi-channel convolutional operations onto physical-layer transformations poses unique challenges, including how to jointly exploit spatial, frequency, and RIS-assisted degrees of freedom (DoFs) to emulate convolution kernels.
In this paper, we propose AirCNN, a novel WPNN framework that emulates 2D CNNs through joint optimization of multi-RIS phase shifts, transmitter precoders, and receiver combiners. Existing methods in [9, 10, 13, 4, 14] cannot be directly extended to realize 2D CNNs. Emulating 2D CNNs poses additional challenges, including incompatibility with traditional data architectures, transceiver design constraints, and transmission protocol limitations. To address these challenges, we investigate two types of 2D CNNs: the classic 2D convolution (Conv2d) and depthwise separable convolution (ConvSD). We propose corresponding transmission architectures and protocols for both RIS-assisted multiple-input single-output (MISO) and multiple-input multiple-output (MIMO) systems. A comprehensive comparison between these two RIS-aided systems for physical CNN realization is conducted, focusing on both performance and communication overhead. Simulation results demonstrate that the proposed methods achieve satisfactory classification accuracy. Moreover, it is shown that multi-RIS setups significantly outperform single-RIS configurations, particularly in line-of-sight (LoS)-dominated wireless channels.
@@ II System model
Fig. 1(a) depicts a conventional CNN architecture, which comprises three modules: initial layers, a middle layer, and final layers. The initial and final layers may contain a large number of neural network layers depending on the specific application, while the middle layer corresponds to the CNN. This paper aims to emulate the CNN layer using wireless hardware to achieve a similar function as shown in Fig. 1(b). In the analog-based CNN architecture, the middle layer consists of precoders at the transmitter, multiple RISs deployed over the air, and combiners at the receiver. We assume that the transmitter and receiver are equipped with $N_{\rm t}$ and $N_{\rm r}$ antennas, respectively. Furthermore, $L$ RISs are deployed, each comprising $M/L$ reflecting elements, where $M$ is the total number of reflecting elements. Let ${{\mathbf{\bar{H}}}_{i}}\in{{\mathbb{C}}^{\left({M/L}\right)\times{N_{\text{t}}}}}$ and ${{{\mathbf{\hat{H}}}}_{i}}\in{{\mathbb{C}}^{{N_{r}}\times\left({M/L}\right)}}$ denote the complex equivalent baseband channel matrices from the transmitter to RIS $i$ and from RIS $i$ to the receiver, respectively. The phase-shift matrix of RIS $i$ is denoted by ${{{\bf{\Theta}}_{i}}}={\rm{diag}}\left({{e^{j{\theta_{i,1}}}},{e^{j{\theta_{i,2}}}},\ldots,{e^{j{\theta_{i,M/L}}}}}\right)$, with $\theta_{i,m}$ denoting the $m$-th phase shift.
@@
A key challenge in AirCNN is the joint design of the transmit precoder, receive combiner, and RIS phase-shift matrices to accurately emulate a given digital convolution kernel $\bf\bar{W}$. To clarify the formulation, we consider a case with multiple kernels under single-channel input and single-channel output. Thus, $\bf\bar{W}$ is a matrix, where each row represents one kernel. This imitation problem can be formulated as [^1] The second term in (1a) explicitly characterizes the effect of receiver noise amplified by the post-processing matrix.
@@
$\qquad\left|{{{\left[{\mathbf{\Theta}}\right]}_{{}_{i,i}}}}\right|=1,{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}i=1,\ldots,M/L,$ (1c)
where ${\bf F}_{1}$ and ${\bf F}_{2}$ denote the transmit precoder and receive combiner, respectively, $P_{\max}$ is the transmit power budget, and $\bf n$ denotes additive white Gaussian noise, following ${\mathbf{n}}\sim{\cal CN}\left({{\bf 0},{\sigma^{2}}{\bf I}}\right)$. The end-to-end channel matrix $\bf H$ is modeled as ${\bf{H}}=\sum\limits_{i=1}^{L}{{{{\bf{\hat{H}}}}_{i}}{{\bf{\Theta}}_{i}}{{{\bf{\bar{H}}}}_{i}}}\in{{\mathbb{C}}^{{N_{r}}\times{N_{\text{t}}}}}$. It is worth noting that problem (1) only characterizes the kernel emulation problem for a single input–output channel pair, where multiple convolution kernels are stacked in $\bf\bar{W}$. Extensions to practical CNN settings with multiple input and output channels are realized by exploiting time, frequency, and spatial degrees of freedom, as detailed in the subsequent Conv2d and ConvSD MISO/MIMO architectures.
The objective of the proposed AirCNN framework is to perform image classification at the receiver, and all trainable components are optimized in an end-to-end manner. Specifically, a standard cross-entropy loss is adopted to train the proposed AirCNN architecture, which can be expressed as
${{\cal L}_{{\text{loss}}}}\left({{\mathbf{\Omega}},{{\mathbf{F}}_{1}},{{\mathbf{F}}_{2}},{\mathbf{\Theta}}}\right)=-\sum\limits_{i=1}^{C}{{p_{i}}\log\left({{{\hat{p}}_{i}}}\right)},$ (2)
where $\mathbf{\Omega}$} denotes the digital neural network parameters, $C$ denotes the number of classes, and $p_{i}$}} and ${\hat{p}}_{i}$}} represent the true label and predicted probability of the $i$th class, respectively.
III Conv2d Realization via OTA Computation
@@ III-A Conv2d MISO
where ${{\mathbf{F}}_{1,i,t}}\in{{\mathbb{C}}^{{N_{{\text{k}}}^{2}}\times{N_{{\text{k}}}^{2}}}}$ and $f_{2,i,t}$$\in{\mathbb{C}}$ denote the $i$-th OFDM carrier-based precoder at the transmitter and the amplification coefficient at the receiver, respectively. In addition, ${{\mathbf{h}}_{t}^{H}}=\sum\limits_{l=1}^{L}{{\mathbf{\hat{h}}}_{l}^{H}{{\mathbf{\Theta}}_{l,t}}{{{\mathbf{\bar{H}}}}_{l}}}$, where ${\bf{\Theta}}_{l,t}$} represents the $l$th RIS phase shift matrix at time slot $t$.
@@ III-B Conv2d MIMO
Remark: In AirCNN, computation is carried out by physical signal propagation rather than sequential digital multiply–accumulate (MAC) operations. Consequently, the dominant complexity does not stem from digital computations, which is also a key motivation for studying WPNNs, but from communication-related overhead, including OTA transmission resources and control signaling.
The communication-related overheads for Conv2d MISO and Conv2d MIMO are summarized in Table I. In the table, $T_{\rm s}$ denotes the number of transmission slots, $T_{\rm r}$ the number of RIS adjustments, $T_{\rm o}$ the number of OFDM carriers, $T_{\rm p}$ the number of precoder adjustments, and $T_{\rm c}$ the number of combiner adjustments.
@@
IV ConvSD Realization via OTA Computation
@@ IV-B ConvSD MIMO
which is similar to (5), except that only a single precoder and a single combiner are required in this case. Then, the receiver applies $C_{\rm out}$ filters, each with $C_{\rm in}$ kernels of size $1\times 1$, to emulate the pointwise convolution process. We have
@@
A communication-related overhead comparison between the parameters of ConvSD MISO and ConvSD MIMO architectures is presented in Table II.
@@ V Numerical results
In this section, we present numerical results to evaluate the image classification accuracy achieved by the proposed schemes, based on the Fashion MNIST dataset. The initial layers consist of one convolutional (Conv) layer, one real-to-complex (R2C) layer, one batch normalization (BN) layer, and one max pool layer, while the final layers have the similar modules, but with an additional complex-to-real (C2R) layer, as the initial layers. We adopt the Adam optimizer with an initial learning rate of $5\times 10^{-4}$, which is decayed by a factor of 0.8 using a multiplicative learning-rate scheduler. The training batch size is fixed to 64. During both training and testing, a new independent channel realization is drawn for each transmitted image, ensuring evaluation under fast-fading conditions. In addition, the Rician fading channel with Rician factor $K$ is considered for both MIMO and MISO systems. Unless otherwise specified, we set $C_{\rm in}=32$, $C_{\rm out}=64$, $N_{\rm wi}=N_{\rm hi}=14$, $N_{\rm k}=3$, $K=3~{\rm dB}$, $N_{t}=9$, $N_{r}=32$ for ConvSD MIMO, $N_{r}=64$ for Conv2d MIMO, $P_{\rm max}=10~{\rm dB}$, and $\sigma^{2}=1$.
@@
In Fig. 4, we study Rician factor $K$ versus classification accuracy for $L=1$, $P_{\rm max}=10~{\rm dB}$, and $M=50$. As $K$ increases, the channel gain improves, resulting in less signal distortion and enhanced classification accuracy. Furthermore, when $K$ is below $10~{\rm dB}$, the ConvSD MISO scheme outperforms the ConvSD MIMO scheme, however, as $K$ increases, ConvSD MIMO scheme eventually surpasses the ConvSD MISO scheme. This behavior is consistent with the explanation provided for Fig. 3. It is also noteworthy that further increase in $K$ does not always lead to continuous improvements in classification accuracy. Beyond a certain threshold, increasing $K$ even degrades the performance, which is discussed in Fig. 6.
@@
In Fig. 6, we study the classification accuracy of the ConvSD MIMO scheme versus $K$ under $P_{\rm max}=10~{\rm dB}$ and $M=50$. It can be observed that for $L=1$, the classification accuracy initially increases but eventually decreases as $K$ continues to grow. This can be explained as follows. When $K$ is below $15~{\rm dB}$, the channel is dominated by non-line-of-sight (NLoS), resulting in a high channel rank and abundant DoFs for neural network modification. Thus, increasing $K$ improves classification accuracy by enhancing the channel gain. However, when $K\geq 15~{\rm dB}$, the channel becomes dominated by LoS components. Although the channel gain remains high, the channel rank tends to decrease toward one, limiting the available DoFs, and thereby, degrading the performance. Moreover, we observe that for larger values of $L$, the system performance remains robust even as $K$ increases. This is because a larger $L$ yields a higher effective channel rank, enhancing the DoFs available for end-to-end training.
@@ VI Conclusion
In this paper, we studied RIS-aided MISO and MIMO systems for engineering the ambient wireless channel to implement CNNs via OTA computation. By jointly training the precoder, combiner, and RIS phase-shift matrices, the digital convolutional operation can be effectively emulated using physical neural networks. We investigated two types of CNNs, namely Conv2d and ConvSD, and proposed two transceiver architectures: RIS-aided MISO and RIS-aided MIMO. A comprehensive comparison between these two architectures for physical CNN realization was conducted, highlighting the trade-offs between performance gains and communication overhead. Simulation results demonstrated that the proposed architectures can achieve satisfactory classification accuracy while enabling OTA computation. Furthermore, it was shown that multi-RIS deployments significantly outperform single-RIS when LoS propagation dominates the wireless channel.